Consider a metal bar with dimensions of l=3.54 cm, w=5.4 cm, and h=10.1 cm

1
Expert's answer
2022-01-19T22:16:01-0500

Solution


Eq. mass of Na2​CO3​=2
Mol.mass​=2
106​=53250 mL of the sodium carbonate solution contains =1.325 g1000 mL of the sodium carbonate solution contains=250
1.325 g​×1000=5.300Normality of Na2​CO3​ solution =Eq.mass
Strength(g/L)​=53
5.30​=10
1​NApplying (Na2​CO3​)N1​V1​​≡(H2​SO4​)N2​V2​​10
1​×25=N2​×20N2​=10×20
25​=8
1​Applying (Before dilution)NB​VB​​≡(After dilution)NA​VA​​8
1​×450=12
1​×VA​VA​=8
450×12​=674 mLWater to be added for dilution =(675−450)=225 mL.